OpenStudy (osanseviero):

In this formula, what is r? I think it is the column of the pascal´s triangle I want, but I am not sure if the first column is column 1 or column 0

OpenStudy (osanseviero):

\[\left(\begin{matrix}n \\ r\end{matrix}\right)a ^{n-r}b ^{r}\]

OpenStudy (osanseviero):

n is the exponent...but r...i dont know

OpenStudy (amistre64):

r is the position of the term in the row

OpenStudy (osanseviero):

but is the first position (always 1) counted? or it counts as term 1?

OpenStudy (osanseviero):

as term 0*

OpenStudy (osanseviero):

For example, if I want the middle term of (a+b)^12, I know I want the sixth term. r should be 5 or 6?

OpenStudy (amistre64):

(a+b)^n, pick the rth term is what im thinking

OpenStudy (amistre64):

(a+b)^3 = a^3 + 3a^2b + 3ab^2 + b^3 hmmm

OpenStudy (amistre64):

3c0 3c1 3c2 3c3 are the coeefs that correspond to ncr

OpenStudy (osanseviero):

ok...then a^3 is term 0

OpenStudy (osanseviero):

so, r could be the term I want -1...thanks

OpenStudy (amistre64):

sounds good :)